OpenStudy (anonymous):

Which choice is equivalent to the fraction below? (Hint: Rationalize the denominator and simplify.) 2 divided by the square root x + square root x+2

OpenStudy (lgbasallote):

do you have a picture or something?

OpenStudy (anonymous):

no it won't let me

OpenStudy (lgbasallote):

ugh..i cant visualize it sorry...i'll call @ParthKohli though

Parth (parthkohli):

\(\color{Black}{\Rightarrow \Large {2 \over \sqrt{x} + \sqrt{x+ 2} } }\) You just multiply both numerator and denominator by \(\sqrt{x} + \sqrt{x + 2}\)